# Beyond Freeze-Out: A Novel Freeze-in Mechanism for Dark Matter via Supercooled Phase Transitions

### v1

@@ Abstract
We investigate a novel freeze-in mechanism for Weakly Interacting Massive Particles (WIMPs), facilitated by a supercooled first-order phase transition (FOPT) in the early universe. Unlike the conventional freeze-out and freeze-in scenarios, this mechanism allows WIMPs to acquire their relic abundance through a non-equilibrium production process following a rapid entropy injection. We explore multiple dark matter (DM) candidates, including vector, fermionic, and scalar-mediated models, and identify fermionic DM with a pseudoscalar mediator as the most viable candidate for this mechanism. The study demonstrates that FOPT dilutes preexisting DM density and simultaneously induces a rapid mass increase, preventing thermal re-equilibration and enabling DM production via freeze-in. Additionally, we analyze the gravitational wave (GW) signals associated with FOPT, identifying parameter regions detectable by future GW observatories such as LISA and UDECIGO. This framework offers a compelling alternative to traditional WIMP and Feebly Interacting Massive Particle (FIMP) scenarios, providing new avenues for DM model building and experimental verification.
@@ 1 Introduction
The role of FOPT and its impact on DM phenomenology is very important[63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. Recently, however, a novel variation of the freeze-in mechanism has been proposed for WIMPs, where weak or moderate interaction strengths suffice due to the influence of a supercooled FOPT in the early universe[97](Freeze-in of WIMP). This scenario opens a new path for DM model building, enabling WIMPs to remain a viable candidate while resolving challenges associated with traditional freeze-out. It provides a robust framework for exploring physics beyond the SM, including connections to FOPT and GW signals. This new mechanism suggests WIMPs can also freeze-in under certain conditions:
@@
The paper is organised as follows. In section. 2, we extend SM by several different models and DM candidates. In section. 3, we examine the conditions of the WIMP freeze-in scenario and present the numerical analyses of the models in light of supercooled FOPT scenario and study production of GWs in section 4, respectively. Finally, the conclusions present in section. 5.
@@ 3 Freeze-in of WIMP
    In the above equation, $g_{*}\approx 106.75$ and for all models presented in the paper, $\Gamma\gg H$, where $\Gamma$ is the decay width of scalars and $H$ is the Hubble constant during the FOPT. The $V_{\Lambda}$ quantity is also defined as follows[63]:
@@
    where $v_{2}$ is the vacuum expectation value of the scalar mediator $S$ and $\beta_{\lambda_{S}}$ is the one-loop beta function for coupling of scalar mediator $\lambda_{S}$. In the following, we examine this condition for all the presented models.
@@
      Refer to caption     Refer to caption
                 (a)                                                                     (b)
      Figure 3: Temperature variations with respect to the independent parameters of the model 2.4. The mixing angle is assumed $\theta=1^{\circ}$. In (a) $v_{2}=200$ GeV, in (b) $M_{H_{2}}=5500$ GeV and in (c) $M_{H_{2}}=5500$ GeV and $v_{2}=200$ GeV is considered.
@@
      This medel includes two independent parameters $M_{V}$ and $g_{v}$. According to equation 2.7, $M_{V}$ must always be greater than 240 GeV. Figure 1 shows the behavior of critical temperature($T_{1}$) and reheating temperature($T_{2}$) for different couplings and masses. As can be seen, only for a very small portion of the parameter space(240 GeV$<M_{V}<$ 500 GeV), $T_{2}$ is larger than $T_{1}$. But to establish the mechanism of Freeze-in of WIMP, we must assume that $T_{2}\gg T_{1}$, which is not the case here, and both temperatures are very close and $T_{2}/T_{1}\sim 1$. Therefore, this model cannot be included in the framework of the presented mechanism.
@@
      Since the model 2.4 with scalar mediator is very limited and it has been publicly ruled out by direct discovery experiments[100], we examine Model 2.4 with pseudoscalar mediator. By choosing the mediator in the model 2.4 to be a pseudoscalar[100], the elastic scattering cross section for DM-nucleon is velocity suppressed and so the theory easily evades the direct detection bounds from XENONnT and LZ[104, 105]. There are five independent parameters $M_{H_{2}}$,$M_{\psi}$,$\theta$, $g_{d}$ and $\nu_{2}$ in this model. An overall result of the Higgs signal strength measured by ATLAS and CMS will restrict the Higgs mixing angle to values smaller than $\sin\theta\leq 0.12$ $(\theta\leq 6.9^{\circ})$[106]. Therefore, in the rest of the paper we consider the mixing angle$(\theta=1^{\circ})$. Also, the mass of DM($M_{\psi}$) will not have any effect on determining the temperatures $T_{1}$ and $T_{2}$. In Figure 3(a), temperature changes are plotted against the mass of $H_{2}$ for constant $g_{d}$ and $\nu_{2}$. As can be seen, as the $M_{H_{2}}$ decreases, $T_{2}$ will become much larger than $T_{1}$, and $\frac{T_{2}}{T_{1}}\gg 1$. In Figure 3(b), temperature changes are plotted against the $\nu_{2}$ for constant $g_{d}$ and $M_{H_{2}}$. As can be seen, as the $\nu_{2}$ decreases, $T_{2}$ will become much larger than $T_{1}$. In Figure 3(c), temperature changes are plotted against the $g_{d}$ for constant $\nu_{2}$ and $M_{H_{2}}$. As can be seen, $T_{2}$ is always greater than $T_{1}$, but their changes are constant with $g_{d}$. As a result, the two determining parameters for examining temperatures are $M_{H_{2}}$ and $\nu_{2}$. Therefore, this model has the initial condition $T_{2}\gg T_{1}$ for the mechanism of freeze-in of WIMP and allows for the examination of this mechanism. It can also be protected from strong constraints of direct detection.
      In Figure 4, the allowed parameter space is also plotted in consistent with the constraint of $T_{2}>T_{1}$ for independent parameters of model. A summary of the analysis performed in this section is provided in Table 3.
So according the above points, we conclude that only fermionic dark matter with a pseudo scalar mediator satisfies the condition of $T_{2}\gg T_{1}$. The supercooled FOPT enables freeze-in WIMP production specifically for it due to a combination of unique conditions and constraints that are satisfied only in this scenario. Here’s a detailed explanation:
  1. 1.
    Entropy Injection and Dilution of Preexisting DM: A supercooled FOPT releases a significant amount of entropy into the universe. This entropy injection dilutes any preexisting DM density to nearly zero. This is a crucial condition for the freeze-in mechanism, as it ensures that the DM production starts from a negligible initial abundance. For fermionic DM with a pseudoscalar mediator, this dilution is particularly effective because the pseudoscalar mediator allows for a specific type of interaction that can maintain the DM out of thermal equilibrium after the phase transition.
  2. 2.
    Mass Increase and Suppression of Thermal Equilibration: During the FOPT, the mass of the DM particles increases significantly. This large mass-to-temperature ratio prevents the DM particles from re-equilibrating with the thermal bath after the phase transition. In the case of fermionic DM with a pseudoscalar mediator, the mass increase is such that the DM particles remain out of equilibrium, which is essential for the freeze-in mechanism to operate. The pseudoscalar mediator facilitates this by providing a coupling that is weak enough to prevent re-equilibration but strong enough to allow for gradual DM production.
@@
Fermionic DM with Pseudoscalar Mediator Possible
@@
    In this scenario, for $T_{2}>T_{dec}$ , the plasma thermalizes again, and the usual freeze-out mechanism yields the relic abundance[101, 63]. $T_{dec}\sim M_{DM}/25$ is temperature of freeze-out. So we are looking for points in the parameter space where $T_{2}<T_{dec}$. Consequently, the evolution of Boltzmann equation begins at $z_{2}=M_{DM}/T_{2}\gg 1$ where the equation is as follows:
    $\frac{dY_{X}}{dz}=-\sqrt{\frac{\pi g^{2}_{*,s}}{45g_{*}}}\frac{M_{Pl}M_{X}}{z^{2}}\langle\sigma\nu_{rel}\rangle(Y_{X}^{2}-Y_{eq}^{2}),$ (3.3)
    where $Y_{X}=n_{X}/s$ is the yield of $X$ with $s$ being the entropy density, $z=M_{X}/T$, $M_{Pl}=1.22\times 10^{19}$ GeV is the Planck scale, $g_{*,s}$ and $g_{*}$ are the numbers the relativistic degrees of freedom for entropy and energy, respectively. $\sigma\nu_{rel}$ is the thermal average of the annihilation cross section multiplying the relative velocity $\nu_{rel}$ under the Maxwell-Boltzmann distribution, and $K_{i}(z)$ is the $i$-th modified Bessel function. Finally, the allowable parameter space according to the Freeze-in WIMP mechanism is plotted in Figure 5. These points are in consistent with the observed relic density of DM[103]. The conditions of $T_{2}\gg T_{1}$ and $T_{2}<T_{dec}$ hold for all points. For these points, $z_{2}$ is between 26 and 32. Of course, we know that $M_{\psi}$ is the mass of DM.
Figure 4: Allowed parameter space in agreement with the condition of $T_{2}>T_{1}$ for the model 2.4 with pseudoscalar mediator . The mixing angle is assumed $\theta=1^{\circ}$. Our scan is done with 4000 GeV$<M_{H_{2}}<$ 6000 GeV, 100 GeV$<\nu_{2}<$ 500 GeV, and the dark coupling in the range 0.01 $<g_{d}<$ 2.
Figure 5: The parameter space that realizes the WIMP freeze-in scenario for model with pseudoscalar mediator. These points are in consistent with the observed relic density of DM, $T_{2}\gg T_{1}$ and $T_{2}<T_{dec}$. The mixing angle is assumed $\theta=1^{\circ}$. Our scan is done with 7 TeV$<M_{\psi}<$ 10 TeV, 100 GeV$<\nu_{2}<$ 200 GeV, 4500 GeV$<M_{H_{2}}<$ 6000 GeV and the dark coupling in the range 0.01 $<g_{d}<$ 2.
4 Gravitational wave signals
@@
The total thermal effective potential is as follows[53]:
@@
We have dropped the Colman-Wienberg zero-temperature correction since at high temperature approximation only the thermal corrections are dominant. In eq.4.3 the parameters $g_{1}$ and $g_{2}$ are respectively the $SU(2)_{L}$ and $U(1)_{Y}$ SM couplings and $g_{t}$ is the top quark coupling.
We use the potential 4.1 to calculate GWs from the model. In the case of significant supercooling, the GW signal arises from bubble collisions[101]. To calculate the GW spectra, we need the phase transition strength $\alpha$, inverse phase transition duration $\beta/H_{*}$, $v_{\omega}$ and $T_{*}(T_{2})$. The bubble collision contribution to the spectrum of GW in the our scenario is given by[109, 110]:
$\Omega_{coll}(f)h^{2}=1.67\times 10^{-5}\Bigl{(}\frac{T_{2}}{T_{\Lambda}}\Bigl{)}^{4/3}\Bigl{(}\frac{\beta}{H_{\ast}}\Bigl{)}^{-2}\Bigl{(}\frac{\kappa\alpha}{1+\alpha}\Bigl{)}^{2}\Bigl{(}\frac{g_{\ast}}{100}\Bigl{)}^{-\frac{1}{3}}\Bigl{(}\frac{0.11v_{\omega}^{3}}{0.42+v_{\omega}^{2}}\Bigl{)}S_{coll},$ (4.5)
where $S_{coll}$ parametrises the spectral shape and is given by
$S_{coll}=\frac{3.8(f/f_{coll})^{2.8}}{2.8(f/f_{coll})^{3.8}+1},$ (4.6)
where
$f_{coll}=1.65\times 10^{-5}\Bigl{(}\frac{T_{2}}{T_{\Lambda}}\Bigl{)}^{1/3}\Bigl{(}\frac{0.62}{\upsilon_{\omega}^{2}-0.1\upsilon_{\omega}+1.8}\Bigl{)}\Bigl{(}\frac{\beta}{H_{\ast}}\Bigl{)}\Bigl{(}\frac{T_{\Lambda}}{100}\Bigl{)}\Bigl{(}\frac{g_{\ast}}{100}\Bigl{)}^{1/6}Hz.$ (4.7)
The $\alpha$ and $\beta/H_{*}$ parameters are as follows[111]:
$\alpha=\frac{\Delta\Bigl{(}V_{eff}-T\frac{\partial V_{eff}}{\partial T}\Bigl{)}\bigg{|}_{T_{*}}}{\rho_{\ast}},$ (4.8)
$\frac{\beta}{H_{\ast}}=T_{*}\frac{d}{dT}\Bigl{(}\frac{S_{3}(T)}{T}\Bigl{)}\bigg{|}_{T_{*}},$ (4.9)
where $\rho_{\ast}$ is
$\rho_{\ast}=\frac{\pi^{2}g_{\ast}}{30}T_{N}^{4}.$ (4.10)
The velocity of the bubble wall($v_{\omega}$) which is anticipated to be close to 1 for the strong transitions[112] and $g_{*}\sim 100$ is the effective number of relativistic degrees of freedom in the thermal plasma. Figure 6 shows the gravitational wave spectrum for the parameters given in table 4. As can be seen, the peak of this spectrum will fall within the range of the LISA and UDECIGO experiments[111, 113].
$M_{\psi}$ (GeV) $g_{d}$ $\nu_{2}$ (GeV) $M_{H_{2}}$ (GeV) $T_{2}$ (GeV) $T_{1}$ (GeV) $\alpha$ $\beta/H_{\ast}$
1 8573 1.26 164.57 5198.14 281.42 1.49 $4.05\times 10^{8}$ 10
2 8390 0.34 140.96 5297.67 286.09 1.5 $2.27\times 10^{9}$ 100
Table 4: Input and output parameters for two benchmark points.
Figure 6: GW spectrum for benchmark points of the Table 4.
5 Conclusion
In the paper, we explored a generalized freeze-in mechanism for DM production, leveraging the effects of a supercooled FOPT. We demonstrated that the entropy injection and mass increase associated with FOPT create conditions that allow WIMPs to undergo freeze-in rather than the conventional freeze-out mechanism. Our study considered multiple DM candidates, including vector, fermionic, and scalar-mediated DM models, identifying scenarios where freeze-in is a viable production mechanism.
Among the models examined, we found that a fermionic DM scenario with a pseudoscalar mediator best satisfies the required conditions for WIMP freeze-in, particularly in maintaining out-of-equilibrium dynamics post-FOPT. This model provides a unique framework where DM relic abundance can be generated with moderate couplings, making it distinct from traditional freeze-in models that typically involve extremely feeble interactions.
Additionally, we analyzed the GW signals associated with the supercooled FOPT, identifying parameter regions where the predicted GW spectra fall within the sensitivity ranges of future experiments such as LISA and UDECIGO. These findings highlight the potential for upcoming GW observations to probe new physics related to DM production.
Overall, this study establishes WIMP freeze-in as a robust and unifying framework for DM model building, offering new perspectives on DM candidates beyond conventional thermal histories. Future research may focus on refining parameter constraints through collider and GW observations, further solidifying the role of FOPT in shaping the early universe’s DM dynamics.

### v2

# Limitations of Freeze-in WIMP Dark Matter from Supercooled Phase Transitions

@@ Abstract
We revisit the possibility of producing Weakly Interacting Massive Particle (WIMP) dark matter via a freeze-in mechanism triggered by a supercooled first-order phase transition (FOPT) in the early universe. Unlike traditional freeze-out and FIMP scenarios, this mechanism relies on a rapid entropy injection that dilutes the preexisting dark matter abundance and prevents re-equilibration due to a sudden mass increase. In this study, we systematically examine a variety of single-component dark matter models—including vector, fermionic, and scalar-mediated candidates—to assess whether they can satisfy the key cosmological condition $T_{2}\gg T_{1}$, required for successful WIMP freeze-in after FOPT. Contrary to earlier results, our revised analysis finds that none of the models fulfill this condition across viable parameter spaces. We confirm, however, that the scalar dark matter model analyzed in Ref. [1] is the only known viable single-component model that fulfills $T_{2}\gg T_{1}$ and enables WIMP freeze-in via this mechanism. These findings place important constraints on model-building efforts and suggest that successful freeze-in after FOPT may require multi-component or more complex dark sectors beyond the scope of minimal models.
@@ 1 Introduction
The role of FOPT and its impact on DM phenomenology is very important[64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98]. Recently, however, a novel variation of the freeze-in mechanism has been proposed for WIMPs, where weak or moderate interaction strengths suffice due to the influence of a supercooled FOPT in the early universe[1](Freeze-in of WIMP). This scenario opens a new path for DM model building, enabling WIMPs to remain a viable candidate while resolving challenges associated with traditional freeze-out. It provides a robust framework for exploring physics beyond the SM, including connections to FOPT and GW signals. This new mechanism suggests WIMPs can also freeze-in under certain conditions:
@@
The paper is organised as follows. In section. 2, we extend SM by several different models and DM candidates. In section. 3, we examine the conditions of the WIMP freeze-in scenario and present the numerical analyses of the models in light of supercooled FOPT scenario. Finally, the conclusions present in section. 4.
@@ 3 Freeze-in of WIMP
    In the above equation, $g_{*}\approx 106.75$ and for all models presented in the paper, $\Gamma\gg H$, where $\Gamma$ is the decay width of scalars and $H$ is the Hubble constant during the FOPT. The $V_{\Lambda}$ is equal to the energy difference between the true and false vacuums during the phase transition. The $V_{\Lambda}$ quantity is also defined as follows[64]:
@@
    where $v_{2}$ is the vacuum expectation value of the scalar mediator $S$ and $\beta_{\lambda_{S}}$ is the one-loop beta function for coupling of scalar mediator $\lambda_{S}$. Of course, equation 3.2 only applies to scale invariant models, i.e. models 2.2 and 2.3. In the following, we examine this condition for all the presented models.
@@
      This medel includes two independent parameters $M_{V}$ and $g_{v}$. According to equation 2.7, $M_{V}$ must always be greater than 240 GeV. Figure 1 shows the behavior of critical temperature($T_{1}$) and reheating temperature($T_{2}$) for different couplings and masses. Details on the calculation of $T_{1}$ are given in the Appendix. As can be seen, only for a very small portion of the parameter space(240 GeV$<M_{V}<$ 500 GeV), $T_{2}$ is larger than $T_{1}$. But to establish the mechanism of Freeze-in of WIMP, we must assume that $T_{2}\gg T_{1}$, which is not the case here, and both temperatures are very close and $T_{2}/T_{1}\sim 1$. Therefore, this model cannot be included in the framework of the presented mechanism.
@@
      Since the model 2.4 with scalar mediator is very limited and it has been publicly ruled out by direct discovery experiments[101], we examine Model 2.4 with pseudoscalar mediator. By choosing the mediator in the model 2.4 to be a pseudoscalar[101], the elastic scattering cross section for DM-nucleon is velocity suppressed and so the theory easily evades the direct detection bounds from XENONnT and LZ[105, 106]. There are five independent parameters $M_{H_{2}}$,$M_{\psi}$,$\theta$, $g_{d}$ and $\nu_{2}$ in this model. An overall result of the Higgs signal strength measured by ATLAS and CMS will restrict the Higgs mixing angle to values smaller than $\sin\theta\leq 0.12$ $(\theta\leq 6.9^{\circ})$[107]. Therefore, in the rest of the paper we consider the mixing angle$(\theta=1^{\circ})$. Also, the mass of DM($M_{\psi}$) will not have any effect on determining the temperatures $T_{1}$ and $T_{2}$.
      Reference [101] contains the equations for calculating critical temperature $T_{1}$. As can be seen, $T_{1}$ is very small, and by calculating the energy difference between true and false vacuums($V_{\Lambda}$), we concluded that $T_{2}$ is of the same order as $T_{1}$, and this model also does not fit into the mechanism of freeze in of WIMP. In fact, the perturbation constraint $\lambda_{S}<4\pi$ causes a limitation in the range of parameter values and leads to a very small critical temperature, which naturally leads to a small temperature $T_{2}$ of the same order as $T_{1}$. Finally, all the models presented in this paper lead to a strong FOPT, but they do not apply to the condition of $T_{2}\gg T_{1}$.
      A summary of the analysis performed in this section is provided in Table 3.
@@
Fermionic DM with Pseudoscalar Mediator Impossible
@@
    In this scenario, for $T_{2}>T_{dec}$ , the plasma thermalizes again, and the usual freeze-out mechanism yields the relic abundance[102, 64]. $T_{dec}\sim M_{DM}/25$ is temperature of freeze-out. So we are looking for points in the parameter space where $T_{2}<T_{dec}$.
4 Conclusion
In this study, we reassessed the freeze-in production of Weakly Interacting Massive Particles (WIMPs) facilitated by a supercooled FOPT in the early universe. The mechanism relies on the sudden injection of entropy and rapid mass gain following FOPT, which dilutes preexisting dark matter and inhibits thermal re-equilibration, thereby enabling freeze-in at moderate couplings.
We analyzed several classes of single-component dark matter models, including vector dark matter with and without scale invariance, fermionic dark matter with U(1) symmetry, and fermionic dark matter with scalar or pseudoscalar mediators. All models were tested against the key requirement $T_{2}\gg T_{1}$, essential for decoupling dark matter from the thermal bath post-FOPT.
Our revised calculations reveal that none of these models satisfy the condition $T_{2}\gg T_{1}$ within viable parameter space. These models thus cannot realize the scenario of freeze-in of WIMP as originally hoped.
The only known single-component model that does meet this condition is the scalar dark matter model previously studied in Ref. [1]. This case remains a unique example where the freeze-in WIMP mechanism can be consistently realized through a supercooled FOPT.
These findings indicate that the minimal freeze-in WIMP framework is highly constrained and may require more complex constructions—such as multi-component dark sectors, richer symmetry structures, or non-standard cosmological histories—to become a broadly viable mechanism. Future work should explore such extended scenarios and their phenomenological signatures, including gravitational wave signals from FOPTs.
Appendix: Calculations related to $T_{1}$ and $\beta_{\lambda_{S}}$
In the high temperature approximation $M_{i}^{2}/T^{2}\ll 1$ for all $i$ with $M_{i}$ being the mass of the particle $i$ in the model, the potential of Model 2.2 is as follows(of course, we know that the potential of the tree level is zero.)[57]:
$V_{eff}^{1-loop}(H_{2},T)=bH_{2}^{4}(\ln\frac{H_{2}^{2}}{\nu^{2}}-\frac{1}{2})+cT^{2}H_{2}^{2},$ (4.1)
where
$b=\frac{1}{64\pi^{2}\nu^{4}}(M_{H_{1}}^{4}+6M_{W}^{4}+3M_{Z}^{4}+3M_{V}^{4}-12M_{t}^{4}),$
$c=\frac{1}{12\nu^{2}}(M_{H_{1}}^{2}+6M_{W}^{2}+3M_{Z}^{2}+3M_{V}^{2}+6M_{t}^{2}).$ (4.2)
The phase transition takes place at the critical temperature $T_{C}$($T_{1}$) at which the finite temperature one-loop effective potential (4.1) has two degenerate minimums at $H_{2}$ = 0 and $H_{2}=\nu_{C}$, i.e.,
$V_{eff}^{1-loop}(0,T_{C})=V_{eff}^{1-loop}(\nu_{C},T_{C})=0,$
$\left.\frac{V_{eff}^{1-loop}(H_{2},T_{C})}{dH_{2}}\right|_{H_{2}=\nu_{C}}=0.$ (4.3)
By simultaneously solving the above equations, the critical temperature is obtained as follows:
$T_{C}=\sqrt{\frac{b}{c}}\nu e^{-\frac{1}{4}}.$ (4.4)
Therefore, the above equation is used to calculate the temperature $T_{1}$ in model2.2. Although the high temperature approximation is used in the calculation of $T_{1}$, the results are no different from examining the full potential form given in reference [100].
Equations 3.1 and 3.2 are used to calculate $T_{2}$. In these equations, the one-loop beta function $\beta_{\lambda_{S}}$ is needed. The beta functions of Model 2.2 are obtained from the following equations, which use the SARAH package[108]:
$(16\pi^{2})\beta_{\lambda_{S}}=20\lambda_{S}^{2}+2\lambda_{SH}^{2}+6g_{v}^{4}-12g_{v}^{2}\lambda_{S},$
$(16\pi^{2})\beta_{\lambda_{SH}}=-\frac{3}{2}g_{1}^{2}\lambda_{SH}-\frac{9}{2}g_{2}^{2}\lambda_{SH}+12\lambda_{SH}\lambda_{H}+8\lambda_{SH}\lambda_{S}+4\lambda_{SH}^{2}+6\lambda_{SH}\lambda_{t}^{2}-6g_{v}^{2}\lambda_{SH},$
$(16\pi^{2})\beta_{\lambda_{H}}=+\frac{3}{8}g_{1}^{4}+\frac{3}{4}g_{1}^{2}g_{2}^{2}+\frac{9}{8}g_{2}^{4}-3g_{1}^{2}\lambda_{H}-9g_{2}^{2}\lambda_{H}+24\lambda_{H}^{2}+\lambda_{SH}^{2}+12\lambda_{H}\lambda_{t}^{2}-6\lambda_{t}^{4},$
$(16\pi^{2})\beta_{g_{v}}=\frac{1}{3}g_{v}^{3}.$ (4.5)
where $\lambda_{t}$ is top Yukawa coupling. The parameters $g_{1}$ and $g_{2}$ are respectively the $SU(2)_{L}$ and $U(1)_{Y}$ SM couplings
Using the results presented above, for Model 2.3 and in the high temperature approximation, the critical temperature is obtained from Equation 4.4 where the coefficients $b$ and $c$ are:
$b=\frac{1}{64\pi^{2}\nu^{4}}(M_{H_{1}}^{4}+6M_{W}^{4}+3M_{Z}^{4}+3M_{V}^{4}-12M_{t}^{4}-4M_{\psi}^{4}),$
$c=\frac{1}{12\nu^{2}}(M_{H_{1}}^{2}+6M_{W}^{2}+3M_{Z}^{2}+3M_{V}^{2}+6M_{t}^{2}+2M_{\psi}^{2}).$ (4.6)
Of course, the results have also been examined with the full potential of the model, as mentioned in reference [54], which does not make much difference.
The beta functions of Model 2.3 are:
$(16\pi^{2})\beta_{\lambda_{S}}=20\lambda_{S}^{2}+4g_{s}^{2}\lambda_{S}-12\lambda_{S}g_{v}^{2}-2g_{s}^{4}+2\lambda_{SH}^{2}+6g_{v}^{4},$
$(16\pi^{2})\beta_{\lambda_{SH}}=-\frac{9}{10}g_{1}^{2}\lambda_{SH}-\frac{45}{10}g_{2}^{2}\lambda_{SH}+12\lambda_{SH}\lambda_{H}+8\lambda_{SH}\lambda_{S}-4\lambda_{SH}^{2}+2\lambda_{SH}g_{s}^{2}-6g_{v}^{2}\lambda_{SH}+6\lambda_{t}^{4},$
$(16\pi^{2})\beta_{\lambda_{H}}=+\frac{27}{200}g_{1}^{4}+\frac{9}{20}g_{1}^{2}g_{2}^{2}+\frac{9}{8}g_{2}^{4}-\frac{9}{5}g_{1}^{2}\lambda_{H}-9g_{2}^{2}\lambda_{H}+24\lambda_{H}^{2}+\lambda_{SH}^{2}+12\lambda_{H}\lambda_{t}^{2}-6\lambda_{t}^{4},$
$(16\pi^{2})\beta_{g_{s}}=\frac{-3}{2}g_{v}^{2}+2g_{s}^{3},$
$(16\pi^{2})\beta_{g_{v}}=g_{v}^{3}.$ (4.7)
Using $\beta_{\lambda_{S}}$ and equations 3.1 and 3.2, the temperature $T_{2}$ can be obtained.
@@ 4 Conclusion
The total thermal effective potential of model 2.4 in the high temperature approximation is as follows[101]:
@@
We have dropped the Colman-Wienberg zero-temperature correction since at high temperature approximation only the thermal corrections are dominant. In eq.4.10 the parameters $g_{1}$ and $g_{2}$ are respectively the $SU(2)_{L}$ and $U(1)_{Y}$ SM couplings and $g_{t}$ is the top quark coupling. Full details and the exact equations required to calculate $T_{1}$ are given in reference [101], which we have used.